Circular No. 7503

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dp IN NGC 1139 S. Beckmann and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Oct. 2.4 and 3.4 UT, both yielding mag ∼ 16.8. The new object is located at α = 2^{h}52^{m}44^{s}.41, δ = −14^{o}31′54′′.7 (equinox 2000.0), which is 34′′.1 west and 8′′.3 south of the nucleus of NGC 1139. A KAIT image of the same field on Sept. 21.4 showed nothing at this position (limiting mag ∼ 18.5).

SATELLITES OF MINOR PLANETS J. L. Margot and M. C. Nolan, National Astronomy and Ionosphere

Center; L. A. M. Benner, S. J. Ostro, R. F. Jurgens, M. A. Slade, and J. D. Giorgini, Jet Propulsion Laboratory (JPL); and D. B. Campbell, Cornell University, report: “Arecibo radar observations (2380 MHz, 12.6 cm) on Sept. 30–Oct. 3 confirm that 2000 DP_{107} is a binary system. Estimates of the diameters based on range extents are 800 and 300 m for the primary and secondary, respectively. Preliminary fits to delay-Doppler data indicate an orbital period of 1.77 days and a semimajor axis of 2.6 km (uncertainties ∼ 10 percent). Based on the above primary diameter, these parameters imply that the density of the primary is ∼ 1.6 g/cm^{3}.”

W. J. Merline, Southwest Research Institute (SRI); L. M. Close, European Southern Observatory; J. C. Shelton, Mount Wilson Observatory; C. Dumas, JPL; F. Menard, Canada-France-Hawaii Telescope (CFHT); and C. R. Chapman and D. C. Slater, SRI, report the discovery of a companion of (90) Antiope on Aug. 10.6 UT from J-, H-, and K′-band direct imaging with the apaptive-optics system of the 10-m W. M. Keck II Telescope on Mauna Kea: “The components differ in brightness by < 0.1 mag and are separated by ∼ 170 km (0′′.12). The orbital plane has a high inclination. The observers continued to acquire observations of the fully resolved components over the next five consecutive nights. Confirming observations of the double were made by Merline, Dumas, and Menard at the CFHT on Sept. 12, 13, and 14. The orbital period is found to be ∼ 16 hr, consistent with the adopted lightcurve period.”

At the recent IAU General Assembly, the satellite of (45) Eugenia (cf. IAUC 7129) was given the permanent designation and name (45) Eugenia I (Petit-Prince).

2000 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green